A satellite orbiting the Earth at 400 km above the surface of the Earth has a 2 m long antenna oriented perpendicular to the Earth's surface. At the equator the Earth's magnetic field is 8×10−5 T and is horizontal. Assuming the orbit to the circular, find emf induced across the ends of the antenna.(Given radius of earth Re ​=6400 Km)


1
Expert's answer
2022-04-19T15:27:24-0400

"v=\sqrt{\frac{gR^2}{R+h}}=7.7~\frac{km}s,"

"\varepsilon=Blv=1.23~V."